OpenStudy (anonymous):

What is the simplified form of \[\frac{ 8x^3 }{ 12y^4 }*\frac{ 20y^6 }{ 6x^8 }\]

OpenStudy (kc_kennylau):

\[\Large\begin{array}{}&\frac{8x^3}{12y^4}\times\frac{20y^6}{6x^8}\\=&\frac{2x^3}{3y^4}\times\frac{10y^6}{3x^8}\\=&\frac{20x^3y^6}{9x^8y^4}\\=&\frac{20y^2}{9x^5}\end{array}\]

OpenStudy (anonymous):

okay i get it now thanks!

OpenStudy (kc_kennylau):

no problem :)